OpenStudy (anonymous):

Dayton is due north of Richmond, and Easton is 16 mi due east of Richmond. Dayton is 32 mi from Easton. How far apart are Dayton and Richmond? Round to the nearest tenth of a mile.

OpenStudy (anonymous):

Ok, by pythogorian theorem 16^2 + x^2=32^2

OpenStudy (anonymous):

x^2=32^2 -16^2 which is.. ?

OpenStudy (anonymous):

768?

OpenStudy (anonymous):

No xD it would be 27.72 when its rounded

OpenStudy (anonymous):

Thanks